OpenStudy (ledah):

Solve by Factoring: x^2 - 2x + 1 = 0

OpenStudy (anonymous):

x^2+2xy+y^2= (x+y)^2 x^2-2xy+y^2= (x-y)^2 i think u will be able to solve it with that knoledge

OpenStudy (anonymous):

its similar to the last one only factor -2 into two parts and b such that a*b=1 a+b=-2

OpenStudy (anonymous):

x^2-2x+1=0 We have a=1,b=-2,c=1 If a+b+c=0 then we have two solutions: First, x1=1, x2=b/a=1 --> x=1 and x=1 (in this case, the same solutions) Or you can say x^2-2x+1 is (x-1)^2-->x=1